John starts his career at 21 years old and expects to retire 44 years later at the age of 65. His first annual salary is $72,000 that will increase at 1.5% per year until he finishes his part-time MBA at 28 years old. With his MBA, John expects salary to increase at 3% per year until retirement. At the end of each year, he deposits 10% of his annual salary into a retirement saving plan that pays 6% interest per year compounded monthly. On the first day of his retirement, John converts his whole retirement saving plan into a registered retirement income fund (RRIF) that earns 8% interest per year compounded quarterly. The RRIF will pay John $Y per quarter, the first payment being paid on the day he buys the RRIF, for 25 years. Find Y. (Show your work without using MS Excel)
Answer #1
Future value(FV) of Cash flow at the end of Year N
FV=(Cash Flow)*(1+i)^(N-t)
t=time (year) of Cash Flow
i=interest rate
Monthly interest rate=(6/12)%= 0.005
Annual effective Rate =((1+0.005)^12)-1= 0.061677812
i=0.061678, N=44
Salary in Year2=72000*1.015= $73,080
Till t=7(Age 28), Salary in year (t+1)=1.015*(Salary in year(t)
After t=7, Salary in year (t+1)=1.03*(salary in year (t)
t A B=0.1*A FV=B*(1.061678^(44-t)
Year Salary Cashflow(Savings) Future Value of Cash flow
1 $72,000 $7,200 $94,411
2 $73,080 $7,308 $90,260
3 $74,176 $7,418 $86,292
4 $75,289 $7,529 $82,498
5 $76,418 $7,642 $78,871
6 $77,564 $7,756 $75,403
7 $78,728 $7,873 $72,088
8 $81,090 $8,109 $69,937
9 $83,522 $8,352 $67,850
10 $86,028 $8,603 $65,826
11 $88,609 $8,861 $63,862
12 $91,267 $9,127 $61,956
13 $94,005 $9,401 $60,108
14 $96,825 $9,683 $58,314
15 $99,730 $9,973 $56,574
16 $102,722 $10,272 $54,886
17 $105,804 $10,580 $53,248
18 $108,978 $10,898 $51,660
19 $112,247 $11,225 $50,118
20 $115,615 $11,561 $48,623
21 $119,083 $11,908 $47,172
22 $122,656 $12,266 $45,765
23 $126,335 $12,634 $44,399
24 $130,125 $13,013 $43,074
25 $134,029 $13,403 $41,789
26 $138,050 $13,805 $40,542
27 $142,191 $14,219 $39,332
28 $146,457 $14,646 $38,159
29 $150,851 $15,085 $37,020
30 $155,376 $15,538 $35,916
31 $160,038 $16,004 $34,844
32 $164,839 $16,484 $33,804
33 $169,784 $16,978 $32,796
34 $174,877 $17,488 $31,817
35 $180,124 $18,012 $30,868
36 $185,527 $18,553 $29,947
37 $191,093 $19,109 $29,053
38 $196,826 $19,683 $28,186
39 $202,731 $20,273 $27,345
40 $208,813 $20,881 $26,529
41 $215,077 $21,508 $25,738
42 $221,530 $22,153 $24,970
43 $228,175 $22,818 $24,225
44 $235,021 $23,502 $23,502
SUM $2,159,580
PV Retirement Savings at age 65 $2,159,580
Rate Quarterly Interest Rate=(8/4)% 2%
Nper Number of quarterly payments 100 (25*4)
Type Payments at beginning of quarters 1
PMT Y=Payment per quarter $49,125.67 (Using PMT function of excel with Rate=2%,Nper=100,Pv=-2159500,Type=1)
ExcelCommand : PMT(2%,100,-2159580,,1)